### v1

@@ Abstract
Smart grids are a fusion of classical power infrastructure and advanced communication networks and smart control, to create a cyber-physical environment that is more efficient and flexible than ever before. This integration causes vulnerabilities that can undermine grid stability as well as reliability. Digital forensics is a fundamental concept of learning and identifying, detecting, and mitigating such security incidents. This paper presents an all-in-one machine learning-based digital forensic framework of smart grid systems deployed on the Cloud. The framework combines the data acquisition at the sensor-level, authenticated communication, scalable cloud storage and automated forensic analytics. The model uses supervised and unsupervised learning algorithms - such as Random Forest, Support Vector Machine, Gradient Boosted Trees and deep neural architectures for anomaly detection, event reconstruction and intrusion analysis in real time. After several simulation and experimental studies on real-time smart-meter data streams, the proposed framework is shown to be very accurate, scalable and resilient to cyber-attacks including data tampering, false-data injection and coordinated control-loop manipulation. The results indicate that cloud services are the best backbone for big-data-driven forensic workflows, which allows energy utilities to achieve a fast situational awareness and intelligent incident response.
Index Terms:
Smart grid, digital forensics, machine learning, anomaly detection, cyber-physical systems, cloud computing, instrumentation analytics, Google Cloud Platform, .
[…]
=-15pt
@@ I Introduction
Rising energy demand, urgent need for decarbonization, increasing penetration of stochastic renewable energy sources (RES), and electrification of transport have necessitated the transformation of the traditional electric grid to an intelligent, adaptive, and sustainable smart grid.[1, 2, 10]. Advancements in sensors have contributed to fusing the grid with a communication network and computational intelligence. The application of these advancements makes the monitoring, control, and optimization of the process possible in real-time across the layers. The aggregation of such a structure is what we know as a cyber-physical system(CPS). [11, 12, 13].
@@
The system complexity and uncertainty at the operational level have significantly increased due to highly flexible loads (EVs) and integration of distributed renewable energy resources (DERs). [3, 14, 15], making traditional model-based control approaches ineffective[16, 4]. Particularly, such centralized architectures often face impediments due to latency, single-point failures, and scalability, reducing their effectiveness for the decentralized structure of the modern smart grids [5, 17]. Therefore, adaptable control architectures with standardized interfaces are warranted. They help resolve the core issue of effective coordination between spatio-temporally dispersed components, since these components may operate under different control regimes. The components may include DER units, microgrids, and EV Charging points.[18, 19, 20].
@@
In addition to spatio-temporal decentrality, the monitoring and control solution for the modern grid needs to be responsive to stochastic events. The monitoring and control solution must be responsive to highly stochastic events such as generation intermittency, load fluctuations, and device-level failures [21, 22]. Moreover, with the increasing fusion of information and communication technologies, the smart grid faces vulnerability vis-à-vis cyber attacks. Cyber attacks such as False Data Injection(FDI) and denial of service(DoS) have a demonstrated capability to undermine control systems and falsify state estimation [23, 24, 25]. Often, these attacks are stealthy and hence compromise the false data detection mechanisms in place, thereby risking the system stability. Lag introduced by the centralized cloud computing platforms is another central issue in the realization of a truly intelligent system. The network delays and resource allocation prove to be limiting the efficacy of otherwise scalable cloud systems. These limitations are especially a problem for time-sensitive grid operations such as voltage control, frequency regulation, and switch protection [26, 27]. Research has been done on edge and fog computing solutions for such problems. These solutions enable localized processing at the edge of the network, thereby improving the response time and resilience of the grid. [28, 29].
@@
Apart from latency, interoperability is also of core concern. While industry standards have been developed to outline guidelines for interfacing and integrating heterogeneous systems [18, 19, 30], ensuring compliance throughout the legacy infrastructure and advanced technologies remains a continuing challenge.
@@
Advancements in artificial intelligence (AI), especially in reinforcement learning and adaptive dynamic programming(ADP), offer a powerful data-driven approach towards finding the solution to complex decision-making problems in the context of increasing uncertainty. These “model-free” approaches have shown promise in direct learning from the environment [6, 7].
@@
Furthermore,agent-based modeling (ABM) and game-theoretic paradigms present robust solutions to represent the behaviors of the grid participants, helping decentralized control and optimization within local and global constraints [8, 9]. However, incorporating AI-based methods within a harmonized hybrid CPS framework, which integrates both cyber and physical layers while ensuring adaptability and security, remains a research gap. Furthermore, limited work has been done on hybrid control schemes focused on combining localized edge intelligence with cloud-based optimization.
@@ II-C1 Optimization Objective
II-D Agen Based and Game Theoretic Modeling
@@ VIII Conclusion
Acknowledgment
The authors would like to sincerely acknowledge the computing lab staff of NFC Institute of Engineering and Technology (IET), Multan, for their technical assistance and support during this research work.

### v2

@@ Abstract
Evolving smart grids require flexible and adaptive control methods. A harmonized hybrid cyber-physical framework, which considers both physical and cyber layers and ensures adaptability, is one of the critical challenges to enable sustainable and scalable smart grids. This paper proposes a three-layer (physical, cyber, control) architecture, with an energy management system as the core of the system. Adaptive Dynamic Programming(ADP) and Artificial Intelligence-based optimization techniques are used for sustainability and scalability. The deployment is considered under two contingencies: Cloud Independent and cloud-assisted. They allow us to test the proposed model under a low-latency localized decision scenario and also under a centralized control scenario. The architecture is simulated on a standard IEEE 33-Bus system, yielding positive results. The proposed framework can ensure grid stability, optimize dispatch, and respond to ever-changing grid dynamics.
@@ I Introduction
Rising energy demand, urgent need for decarbonization, increasing penetration of stochastic renewable energy sources (RES), and electrification of transport have necessitated the transformation of the traditional electric grid to an intelligent, adaptive, and sustainable smart grid.[fang2012smart, amin2005toward, farhangi2010path]. Advancements in sensors have contributed to fusing the grid with a communication network and computational intelligence. The application of these advancements makes the monitoring, control, and optimization of the process possible in real-time across the layers. The aggregation of such a structure is what we know as a cyber-physical system(CPS). [gungor2011smart, mohsenian2010autonomous, momoh2012smart].
@@
The system complexity and uncertainty at the operational level have significantly increased due to highly flexible loads (EVs) and integration of distributed renewable energy resources (DERs). [zhang2011game, chakraborty2016power, yang2019distributed], making traditional model-based control approaches ineffective[liu2009cyber, teixeira2015secure]. Particularly, such centralized architectures often face impediments due to latency, single-point failures, and scalability, reducing their effectiveness for the decentralized structure of the modern smart grids [gao2012cyber, huang2018review]. Therefore, adaptable control architectures with standardized interfaces are warranted. They help resolve the core issue of effective coordination between spatio-temporally dispersed components, since these components may operate under different control regimes. The components may include DER units, microgrids, and EV Charging points.[ieeep2030, ieee1547, yan2013survey].
@@
In addition to spatio-temporal decentrality, the monitoring and control solution for the modern grid needs to be responsive to stochastic events. The monitoring and control solution must be responsive to highly stochastic events such as generation intermittency, load fluctuations, and device-level failures [khanna2016energy, palensky2011demand]. Moreover, with the increasing fusion of information and communication technologies, the smart grid faces vulnerability vis-à-vis cyber attacks. Cyber attacks such as False Data Injection(FDI) and denial of service(DoS) have a demonstrated capability to undermine control systems and falsify state estimation [liu2011false, sridhar2012cyber, rawat2015cybersecurity]. Often, these attacks are stealthy and hence compromise the false data detection mechanisms in place, thereby risking the system stability. Lag introduced by the centralized cloud computing platforms is another central issue in the realization of a truly intelligent system. The network delays and resource allocation prove to be limiting the efficacy of otherwise scalable cloud systems. These limitations are especially a problem for time-sensitive grid operations such as voltage control, frequency regulation, and switch protection [alam2017fog, tao2017edge]. Research has been done on edge and fog computing solutions for such problems. These solutions enable localized processing at the edge of the network, thereby improving the response time and resilience of the grid. [kulkarni2019edge, mohagheghi2010distributed].
@@
Apart from latency, interoperability is also of core concern. While industry standards have been developed to outline guidelines for interfacing and integrating heterogeneous systems [ieeep2030, ieee1547, smartgridinterop], ensuring compliance throughout the legacy infrastructure and advanced technologies remains a continuing challenge.
@@
Advancements in artificial intelligence (AI), especially in reinforcement learning and adaptive dynamic programming(ADP), offer a powerful data-driven approach towards finding the solution to complex decision-making problems in the context of increasing uncertainty. These “model-free” approaches have shown promise in direct learning from the environment [liu2019reinforcement, chen2021deep].
@@
Furthermore,agent-based modeling (ABM) and game-theoretic paradigms present robust solutions to represent the behaviors of the grid participants, helping decentralized control and optimization within local and global constraints [saad2012game, sharma2018game]. However, incorporating AI-based methods within a harmonized hybrid CPS framework, which integrates both cyber and physical layers while ensuring adaptability and security, remains a research gap. Furthermore, limited work has been done on hybrid control schemes focused on combining localized edge intelligence with cloud-based optimization.
@@ II-C1 Optimization Objective
II-D Agent‑Based and Game‑Theoretic Modeling